# Image Generation Models: A Technical History

### 08 Mar 2026 v1

@@ 2 Variational Autoencoders
where $\phi$ and $\theta$ are encoder and decoder networks and $z\in\mathbb{R}^{d}$ where $d<D$. The two networks are trained by minimizing the reconstruction error between $x$ and $\hat{x}$ through the following MSE (Min Squared Error) loss term:
@@
Rumelhart, Hinton, and Williams [2] introduced the idea of compressing the input into a bottleneck representation and reconstructing it back. Vincent et al. [3] showed that corrupting inputs and training the model to reconstruct them produces robust feature representations. Yoshua Bengio and his coauthors [4] generalized this framework, proving that denoising autoencoders can be seen as consistent estimators of the data distribution and proposing the Markov-based training procedure to improve the quality of generative sampling. Variational Autoencoders (VAEs) improve the autoencoders in two ways. First, the autoencoders did not enforce structure on the latent space. On the other hand, VAEs used the Kullback-Leibler (KL) term to force the latent space $z$ to follow a particular distribution, Eq. 3.
@@
As shown in Fig. 1, In early generations of autoencoders a randomly sampled latent code $z$ often decoded to meaningless outputs, VAEs regularize the latent space so that random samples can be reliably mapped back into the data space. VAEs force $z$ to follow a particular probability distribution via prior distribution $p(z)$. As a result, $z$ is no longer irregular with holes and rather a smooth distribution.
@@
Second, VAEs introduced a structured likelihood based training structure which provided a clean mathematical framework for training both the encoder and decoder simultaneously.
@@ 2.1 How VAEs work?
Given the model parameters $\theta$, we are interested in knowing how likely it is to see datapoint $x$, $p_{\theta}(x)$. In order to maximize the likelihood of seeing $x$, we would like to fit it such that the probability assigned to real data points $x$ is as high as possible, Eq. 4.
@@
This integral sums over all possible latent codes $z$ that could have generated $x$. In high-dimensional space, calculating this integral is intractable. Instead, we use a neural network $q_{\phi}(z\mid x)$ (the encoder) that gives a tractable approximation to the distribution of $z$ given $x$. We can rewrite:
@@
Applying Jensen’s inequality, $\log\mathbb{E}[X]\geq\mathbb{E}[\log X]$, results in
@@
Since $p_{\theta}(x,z)=p_{\theta}(x\mid z)\,p(z)$, ELBO can be rewritten as
@@
The reconstruction term forces the decoder $p_{\theta}(x\mid z)$ to explain $x$ well when $z$ is sampled from the encoder $q_{\phi}(z\mid x)$. The regularizer $\mathrm{KL}(q_{\phi}(z\mid x)\,\|\,p(z))$ pushes the approximate posterior closer to the prior $p(z)$. In practice, the negative ELBO is minimized as the loss function
@@
In implementations, $q_{\phi}(z\mid x)$ and $p_{\theta}(x\mid z)$ are estimated by encoder and decoder neural networks. $z$ is commonly drawn from a Gaussian prior, i.e., $z\sim\mathcal{N}(0,I)$
@@ 2.2 Reparameterization Trick
This approach leads to a much lower variance, and end-to-end backpropagation becomes possible. Intuitively, this means that they are sampling from fixed noise and pushing it through a differentiable function that depends on $\phi$. Almost simultaneously to [5], Rezende et al. [6] published a similar work where they presented stochastic backpropagation as general rules for backpropagating through stochastic nodes and derived gradient identities for Gaussians. Both of these papers introduced reparameterization and also the idea of backpropagation through stochastic nodes.
@@ 2.3 KL Collapse: when the model ignores $z$
where $C_{t}$ gradually increases from 0 to $C_{\max}$ over the course of training. Intuitively, early in training when the encoder has very limited capacity, it must focus on transmitting only the most useful coarse factors. As $C_{t}$ grows, the encoder gains more capacity to capture fine-grained details.
@@ 2.4 Blurry Reconstructions in VAEs
Oord et al. [9] introduced PixelCNN for conditional image generation where they predict each pixel from the surrounding pixels in a raster order:
@@
where $x_{<i}$ are earlier pixels in raster order. As shown in Fig. 3 PixelCNNs [9] are capable of producing sharper images since they model a joint distribution.
@@
Gulrajani et al. [11] introduced PixelVAE where they use PixelCNN to model an autoregressive decoder in the decoder section of the VAE. This can help VAEs generate sharper output images.
$p(z\mid x)=\prod_{i=1}^{HW}p(z_{i}\mid x,z_{<i}),\quad z\sim p(z).$ (21)
@@ 2.5 Conditional VAEs
So far we have talked about the generation process in VAEs without much emphasis on controlling the generated output. Sohn et al. [12] introduced Conditional VAE (CVAE), where the generated output is conditioned on the input. Fig. 5 shows a simplified diagram of conditional VAE. Here are a few examples where the condition is represented as input $x$ and a target represented by $y$:
  • input image $x\to$ segmented image $y$
  • painted input $x\to$ realistic photo $y$
Figure 5: Conditional VAE: You can think of $X$ as the condition upon which the output is being generated.
@@
The training and inference steps for a conditional VAEs can be simplified as shown in Box 1.
@@ 2.6.2 Vector Quantized VAE
In Vector Quantized Autoencoder(VQ-VAE), Oord et al. [1] try to enhance VAEs in two ways. First, the encoder network outputs a discrete (quantized) value, named codebook. This reduces the need for the encoder to average over fine details and instead it can learn to assign discrete codes to individual image attributes, resulting in sharper reconstructions with less blurry output images. Second, the prior $p(z)$ is learned and is no longer a fixed Gaussian distribution. As we will see in Section 6, the overall concept of VQ-VAEs play a central role in several successful diffusion models. Fig. 8 shows the overall diagram.
@@
As shown in Box 2, the training happens in two stages. In stage 1, the encoder, codebook, and decoder are trained to find the closest value in the codebook that matches the output of the encoder. In stage 2, the encoder and decoder are frozen and the prior is trained over code indices. An autoregressive prior $p_{\psi}(c)$ (e.g., PixelCNN, Transformer) can be trained on the code sequences:
Training Steps for VQ-VAE Stage 1 1. $z_{e}=\text{Encoder}_{\phi}(x),\quad z_{e}\in\mathbb{R}^{D\times H\times W}$ 2. Vector Quantization (nearest neighbor per location): $k^{*}(u,v)$ $=\arg\min_{k}\left\lVert z_{e}(:,u,v)-e_{k}\right\rVert_{2},$ $z_{q}(:,u,v)$ $=e_{k^{*}(u,v)}\in\mathbb{R}^{D}$ 3. Straight-Through (ST) estimator for backpropagation: $z_{q}^{ST}=z_{e}+(z_{q}-z_{e})_{\text{stop-grad}}$ Forward path: $z_{q}^{ST}=z_{q}$ Backward path: $\tfrac{\partial z_{q}^{ST}}{\partial z_{e}}=I,\quad\tfrac{\partial z_{q}^{ST}}{\partial z_{q}}=0$ 4. $\hat{x}=\text{Decoder}_{\theta}(z_{q}^{ST})$ 5. Loss: $\mathcal{L}=\mathcal{L}_{\text{reconstruction}}(x,\hat{x})+\beta\lVert z_{e}-\mathrm{sg}[z_{q}]\rVert_{2}^{2}$ where $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. 6. Update parameters $\theta,\phi$ via backpropagation, and update embeddings $E$ via exponential moving average (EMA). Stage 2 1. Freeze the encoder/decoder networks and codebook. 2. Let $c\in\{1,\dots,K\}^{H\times W}$ be the grid of code indices $c(u,v)=k^{*}(u,v)$ from Stage 1. 3. Train an autoregressive prior $p_{\psi}(c)$ (e.g., PixelCNN/Transformer) on the index sequences: $p_{\psi}(c)\;=\;\prod_{i=1}^{HW}p_{\psi}\!\big(c_{i}\,\big|\,c_{<i}\big),$ 4. Use cross-entropy loss and update the parameters $\phi$
@@ 2.6.3 Deep Hierarchical VAEs
Vahdat and Kautz [21] introduced Nouveau VAE(NVAE) where they scale up the training into a large number of hierarchical groups and image sizes while keeping the training stable via regularization and reparameterization. Using this approach they are able to generate high-quality images like Fig. 10 on the CelebA dataset [22]
@@
Rewon Child [23] took this one step further and introduced a Very Deep Variational Autoencoder(VDVAE). By making the hierarchy very deep like in Fig. 11, the VAE’s top-down chain of priors acts as autoregressor over the latent space.
@@ 2.7 Conclusion
Although VAEs offered a strong probabilistic foundation for generating high-quality images, they also suffer from several shortcomings. First, as discussed earlier, the ELBO loss is only an approximation of the true likelihood. Moreover, decoders can sometimes ignore the latent space, which can lead to posterior collapse. Lastly, despite the interpretability and controllability advantages of VAEs, they are often difficult to train and optimize. In particular, large gradients can destabilize the training regime. Later, better optimization techniques were utilized in order to make VAE training more stable. As we will see in Section 6 VQ-VAEs made their way through next generations of image generations such as diffusion models and played a crucial role in their success.
@@ 3 Generative Adversarial Networks
Generative Adversarial Networks (GANs) were introduced by Goodfellow et al. [24] only one year after Variational Autoencoders. Although different in approach, the goal was similar, generating high-quality images. Their approach uses two models: a generative model $G$ and a discriminative model $D$. They used a multi-layer perceptron for both discriminator and generator.
@@
As a result the generator and discriminator’s cost functions can be written separately as below.
@@
Early during training, $G$ is weak, and $D$ can easily reject fake samples. As a result, instead of training $G$ to minimize $\log(1-D(G(z)))$, they train $G$ to maximize $\log D(G(z))$. This small change provides a much stronger gradient early in training. Additionally, in training frameworks, the preference is to work with minimization functions; as a result, we can change the objective functions to minimization. Applying these two changes modifies the objective functions as shown below.
Training Algorithm for GANs for $N$ training iterations do for $k$ steps do Sample minibatch of $m$ noise samples $\{z^{(1)},\ldots,z^{(m)}\}$ from noise prior $p_{z}(z)$. Sample minibatch of $m$ examples $\{x^{(1)},\ldots,x^{(m)}\}$ from data generating distribution $p_{\text{data}}(x)$ Update the discriminator by descending its stochastic gradient: $\nabla_{\!\theta_{d}}\frac{1}{m}\sum_{i=1}^{m}\Big[-\log D(x^{(i)})-\log\big(1-D(G(z^{(i)}))\big)\Big]$ end for Sample minibatch of $m$ noise samples $\{z^{(1)},\ldots,z^{(m)}\}$ from noise prior $p_{z}(z)$. Update the generator by descending its stochastic gradient: $\nabla_{\!\theta_{g}}\frac{1}{m}\sum_{i=1}^{m}-\log\big(D(G(z^{(i)}))\big)$ end for
@@
The training process can be simplified as shown in Box 3. Fig. 13 shows some of the sample images drawn from the generator $G$. One of the major advantages of GANs is the fact that the whole network is trained end-to-end using backpropagation. Another advantage of GANs is that the generator network is not directly updated from data example, but only with gradients flowing through the discriminator network. This can help since the discriminator can turn data into differentiable objective for the generator. One major challenge the authors faced was the training instability. For example, the discriminator network $D$ must be in sync with the generator network $G$ during training. If $G$ is trained too much without updating $D$, then $G$ collapses many values of $z$ to the same value of $x$. They remedy this by alternating between $k$ optimization steps for $D$ and one optimization step for $G$.
@@
Building on the success of Convolutional Neural Networks (CNNs) and Generative Adversarial Networks (GANs), Radford et al. [25] introduced Deep Convolutional Generative Adversarial Networks (DCGANs). They introduced a set of architectural changes to the topology of GANs that offered a more stable training. The architectural properties of the later GAN papers can be traced back to this paper. They replaced pooling layers with strided convolutions in the $D$ and fractional-strided/transpose convolutions in the $G$. This tweak allows the model to learn downsampling/upsampling; moreover, it helps with better gradient flow and spatial control. They also utilized batch normalization in intermediate layers of both $G$ and $D$ to stabilize training and prevent early collapse. Moreover, They replaced fully connected layers with stacking deep convolutional layers. They only used fully connected layers in the $G$ for projection to a 4D tensor. Lastly, They used ReLU in the $G$ (except for tanh in the output) and LeakyReLU in $D$ for stronger gradients.
Looking at the results of their work, they were able to generate stable $64\times 64$ images on LSUN Bedrooms, ImageNet-1k crops, and Faces datasets. Their results, as shown in Fig. 14, seem cleaner than the prior GAN results.
@@
They also analyze latent-space arithmetic. They performed vector arithmetic in the latent space:
@@
They then fed $z_{\text{new}}$ to the generator. The result was images that semantically reflected the arithmetic, as shown in Fig. 15.
This was a strong hint that the GANs’ latent space is organized linearly enough that simple vector operations map to interpretable images. Moreover, it shows that GANs are able to learn features that are disentangled.

### 29 Mar 2026 v2

@@ 2 Variational Autoencoders
where $\phi$ and $\theta$ are encoder and decoder networks and $z\in\mathbb{R}^{d}$ where $d<D$. The two networks are trained by minimizing the reconstruction error between $x$ and $\hat{x}$ through the following MSE (Mean Squared Error) loss term:
@@
Rumelhart, Hinton, and Williams [2] introduced the idea of compressing the input into a bottleneck representation and reconstructing it back. Vincent et al. [3] showed that corrupting inputs and training the model to reconstruct them produces robust feature representations. Yoshua Bengio and his coauthors [4] generalized this framework, proving that denoising autoencoders can be seen as consistent estimators of the data distribution and proposing the Markov-based training procedure to improve the quality of generative sampling. Variational Autoencoders (VAEs) improve the autoencoders in two ways. First, the autoencoders did not enforce structure on the latent space. On the other hand, VAEs use the Kullback-Leibler (KL) term to force the latent space $z$ to follow a particular distribution such as Gaussian as shown in Eq. 3.
@@
As shown in Fig. 1 (top), in early generations of autoencoders a randomly sampled latent code $z$ often decoded to meaningless outputs, VAEs regularize the latent space so that random samples can be reliably mapped back into the data space. VAEs force $z$ to follow a particular probability distribution via prior distribution $p(z)$. As a result, $z$ is no longer irregular with holes and rather a smooth distribution as shown in Fig. 1 (bottom).
Second, VAEs introduce a structured likelihood-based training structure that provides a clean mathematical framework for training both the encoder and decoder simultaneously.
@@ 2.1 How VAEs work?
Given the model parameters $\theta$, we are interested in knowing how likely it is to see the real data $x$, $p_{\theta}(x)$. In order to maximize the likelihood of seeing $x$, we would like to fit it such that the probability assigned to real data points $x$ is as high as possible, Eq. 4.
@@
This integral sums over all possible latent codes $z$ that could have generated $x$. In high-dimensional space, calculating this integral is intractable. Instead, we use a neural network $q_{\phi}(z\mid x)$ (the encoder) that gives a tractable approximation of the distribution of $z$ given $x$. We can rewrite:
@@
Applying Jensen’s inequality, $\log\mathbb{E}[X]\geq\mathbb{E}[\log X]$, results in:
@@
Since $p_{\theta}(x,z)=p_{\theta}(x\mid z)\,p(z)$, ELBO can be rewritten as:
@@
The reconstruction term forces the decoder $p_{\theta}(x\mid z)$ to explain $x$ well when $z$ is sampled from the encoder $q_{\phi}(z\mid x)$. The regularizer $\mathrm{KL}(q_{\phi}(z\mid x)\,\|\,p(z))$ pushes the approximate posterior (encoder output) closer to the prior $p(z)$. In practice, the negative ELBO is minimized as the loss function
@@
In implementations, $q_{\phi}(z\mid x)$ and $p_{\theta}(x\mid z)$ are estimated by encoder and decoder neural networks as depicted in Fig. 2. $z$ is commonly drawn from a Gaussian prior, i.e., $z\sim\mathcal{N}(0,I)$
@@ 2.2 Reparameterization Trick
This approach leads to a much lower variance, and end-to-end backpropagation becomes possible. Intuitively, this means that they are sampling from fixed noise and pushing it through a differentiable function that depends on $\phi$. Almost simultaneously to [5], Rezende et al. [6] published a similar work where they presented stochastic backpropagation as a general rule for backpropagating through stochastic nodes and derived gradient identities for Gaussians. Both of these papers introduced reparameterization and also the idea of backpropagation through stochastic nodes.
@@ 2.3 KL Collapse: when the model ignores $z$
where $C_{t}$ gradually increases from 0 to $C_{\max}$ over the course of training. Early in training when the encoder has very limited capacity, it must focus on transmitting only the most useful coarse factors. As $C_{t}$ increases, the encoder gains more capacity to capture fine-grained details.
@@ 2.4 Blurry Reconstructions in VAEs
In order to alleviate this problem, Oord et al. [9] introduced PixelCNN for conditional image generation where they predict each pixel from the surrounding pixels in a raster order:
@@
where $x_{<i}$ are earlier pixels in raster order. As shown in Fig. 3 PixelCNNs are capable of producing sharper images since they model a joint distribution.
@@
Later, Gulrajani et al. [11] introduced PixelVAE where they use PixelCNN to model an autoregressive decoder in the decoder section of the VAE. This can help VAEs generate sharper output images.
$p_{\theta}(x\mid z)=\prod_{i=1}^{HW}p_{\theta}(x_{i}\mid x_{<i},z),\quad z\sim p(z).$ (21)
@@ 2.5 Conditional VAEs
So far we have talked about the generation process in VAEs without much emphasis on controlling the generated output. Sohn et al. [12] introduced Conditional VAE (CVAE), where the generated output is conditioned on the input. Fig. 5 shows a simplified diagram of conditional VAE. Here are two examples where the condition is represented as input $x$ and a target represented by $y$ in the $(x,y)$ pairs: (input image, segmented image) or (painted input, realistic photo).
Figure 5: Conditional VAE: One can think of $X$ as the condition upon which the output is being generated.
@@
The training and inference steps for conditional VAEs can be simplified as shown in Box 1.
@@ 2.6.3 Vector Quantized VAE
In Vector Quantized Autoencoder (VQ-VAE), Oord et al. [1] try to enhance VAEs in two ways. First, the encoder outputs latent vectors that are quantized by selecting entries from a learned codebook. This reduces the need for the encoder to average over fine details and instead it can learn to assign discrete codes to individual image attributes, resulting in sharper reconstructions with less blurry output images. Second, the prior $p(z)$ is learned and is no longer a fixed Gaussian distribution. As we will see in Section 6, the overall concept of VQ-VAEs plays a central role in several successful diffusion models. Fig. 8 shows the overall diagram.
@@
As shown in Box 2, training for VQ-VAE occurs in two stages. In stage 1, the encoder, codebook, and decoder are trained to find the closest value in the codebook that matches the output of the encoder. In stage 2, the encoder and decoder are frozen and the prior is trained over code indices. An autoregressive prior $p_{\psi}(c)$ (e.g., PixelCNN, Transformer) can be trained on the code sequences:
Training Steps for VQ-VAE Stage 1 1. $z_{e}=\text{Encoder}_{\phi}(x),\quad z_{e}\in\mathbb{R}^{D\times H\times W}$ 2. Vector Quantization (nearest neighbor per location): $k^{*}(u,v)$ $=\arg\min_{k}\left\lVert z_{e}(:,u,v)-e_{k}\right\rVert_{2},$ $z_{q}(:,u,v)$ $=e_{k^{*}(u,v)}\in\mathbb{R}^{D}$ 3. Straight-Through (ST) estimator for backpropagation: $z_{q}^{ST}=z_{e}+(z_{q}-z_{e})_{\text{stop-grad}}$ Forward path: $z_{q}^{ST}=z_{q}$ Backward path: $\tfrac{\partial z_{q}^{ST}}{\partial z_{e}}=I,\quad\tfrac{\partial z_{q}^{ST}}{\partial z_{q}}=0$ 4. $\hat{x}=\text{Decoder}_{\theta}(z_{q}^{ST})$ 5. Loss: $\mathcal{L}=\mathcal{L}_{\text{reconstruction}}(x,\hat{x})+\beta\lVert z_{e}-\mathrm{sg}[z_{q}]\rVert_{2}^{2}$ where $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. 6. Update parameters $\theta,\phi$ via backpropagation, and update embeddings $E$ via exponential moving average (EMA). Stage 2 1. Freeze the encoder/decoder networks and codebook. 2. Let $c\in\{1,\dots,K\}^{H\times W}$ be the grid of code indices $c(u,v)=k^{*}(u,v)$ from Stage 1. 3. Train an autoregressive prior $p_{\psi}(c)$ (e.g., PixelCNN/Transformer) on the index sequences: $p_{\psi}(c)\;=\;\prod_{i=1}^{HW}p_{\psi}\!\big(c_{i}\,\big|\,c_{<i}\big),$ 4. Use cross-entropy loss and update the parameters $\psi$.
@@ 2.6.4 Deep Hierarchical VAEs
Later, Vahdat and Kautz [21] introduced Nouveau VAE (NVAE) where they scale up the training into a large number of hierarchical groups and image sizes while keeping the training stable via regularization and reparameterization. Using this approach, they are able to generate high-quality images like Fig. 10 on the CelebA dataset [22]
@@
Rewon Child [23] took this one step further and introduced a Very Deep Variational Autoencoder (VDVAE). By making the hierarchy very deep like in Fig. 11, the VAE’s top-down chain of priors acts as an autoregressor over the latent space.
@@ 2.7 Conclusion
Although VAEs offered a strong probabilistic foundation for generating high-quality images, they also suffer from several shortcomings. First, as discussed earlier, the ELBO loss is only an approximation of the true likelihood. Moreover, decoders can sometimes ignore the latent space, which can lead to posterior collapse. Lastly, despite the interpretability and controllability advantages of VAEs, they are often difficult to train and optimize. In particular, large gradients can destabilize the training regime. Later, better optimization techniques were utilized in order to make VAE training more stable. As we will see in Section 6 VQ-VAEs made their way through next generations of image synthesis such as diffusion models and played a crucial role in their success.
@@ 3 Generative Adversarial Networks
Generative Adversarial Networks (GANs) were introduced by Goodfellow et al. [24] shortly after Variational Autoencoders. Although different in approach, the goal was similar, generating high-quality images. Their approach uses two models: a generative model $G$ and a discriminative model $D$. They used a multi-layer perceptron for both discriminator and generator.
@@
One can write the generator and discriminator’s cost functions separately as in the following:
@@
Early during training, $G$ is weak, and $D$ can easily reject fake samples. As a result, instead of training $G$ to minimize $\log(1-D(G(z)))$, they train $G$ to maximize $\log D(G(z))$. This small change provides a much stronger gradient early in training. Additionally, in training frameworks, the preference is to work with minimization functions; as a result, we can change the objective functions to minimization. Applying these two changes modifies the objective functions as shown in Eq. 3536.
Training Algorithm for GANs for NN training iterations do for $k$ steps do Sample minibatch of $m$ noise samples $\{z^{(1)},\ldots,z^{(m)}\}$ from noise prior $p_{z}(z)$. Sample minibatch of $m$ examples $\{x^{(1)},\ldots,x^{(m)}\}$ from data generating distribution $p_{\text{data}}(x)$ Update the discriminator by descending its stochastic gradient: $\nabla_{\!\theta_{d}}\frac{1}{m}\sum_{i=1}^{m}\Big[-\log D(x^{(i)})-\log\big(1-D(G(z^{(i)}))\big)\Big]$ end for Sample minibatch of $m$ noise samples $\{z^{(1)},\ldots,z^{(m)}\}$ from noise prior $p_{z}(z)$. Update the generator by descending its stochastic gradient: $\nabla_{\!\theta_{g}}\frac{1}{m}\sum_{i=1}^{m}-\log\big(D(G(z^{(i)}))\big)$ end for
@@
One of the major advantages of GANs is the fact that the whole network is trained end-to-end using backpropagation. Another advantage of GANs is that the generator network is not directly updated from data examples, but only with gradients flowing through the discriminator network. This can help since the discriminator can turn data into a differentiable objective for the generator. One major challenge the authors faced was the training instability. For example, the discriminator network $D$ must be in sync with the generator network $G$ during training. If $G$ is trained too much without updating $D$, then $G$ collapses many values of $z$ to the same value of $x$. They remedy this by alternating between $k$ optimization steps for $D$ and one optimization step for $G$. The training process can be simplified as shown in Box 3. Fig. 13 shows some of the sample images drawn from the generator $G$.
@@
Building on the success of Convolutional Neural Networks (CNNs) and GANs, Radford et al. [25] introduced Deep Convolutional Generative Adversarial Networks (DCGANs). They introduced a set of architectural changes to the topology of GANs that offered a more stable training. In fact, the architectural properties of the later GAN papers can be traced back to this paper. They replaced pooling layers with strided convolutions in the $D$ and fractional-strided/transpose convolutions in the $G$. This tweak allows the model to learn downsampling/upsampling; moreover, it helps with better gradient flow and spatial control. They also utilized batch normalization in intermediate layers of both $G$ and $D$ to stabilize training and prevent early collapse. They also replaced fully connected layers with stacking deep convolutional layers. They only used fully connected layers in the $G$ for projection to a 4D tensor. Lastly, They used ReLU in the $G$ (except for tanh in the output) and LeakyReLU in $D$ for stronger gradients.
Looking at the results of their work, they were able to generate stable $64\times 64$ images on the LSUN Bedrooms, ImageNet-1k crops, and Faces datasets. Their results, as shown in Fig. 14, seem cleaner than the prior GAN results.
@@
They also analyze latent-space arithmetic. They performed vector arithmetic in the latent space via Eq. 37:
@@
They then fed $z_{\text{new}}$ to the generator. The result was images that semantically reflected the arithmetic, as shown in Fig. 15. This was a strong hint that the GANs’ latent space is organized linearly enough that simple vector operations map to interpretable images. Moreover, it shows that GANs are able to learn features that are disentangled.